OpenStudy (anonymous):

How do I solve: Given: csc θ = −(17/15), where 270° ≤ θ ≤ 360° and cot β = −(3/4) where 90° ≤ β ≤ 180°. Find the exact value of sin(θ + β).

OpenStudy (callisto):

csc θ = −(17/15) => 1/sinθ = −(17/15) => sinθ = -15/17 opp side= 15, hyp. =17 By Pyth Thm, we get adj. side =8 => cos θ = 8/17 (since it is in quad IV) cot β = −(3/4) => 1/tanβ = −(3/4) => tanβ = -4/3 opp. side =4, adj side =3, By Pyth Thm, hyp. =5 => sinβ = 4/5 cosβ = -3/5 (since it is in quadrant II) sin(θ + β) = sinθ cosβ + cosθsinβ = (-15/17) (-3/5) + ( 8/17)(4/5) Can you do it now?

OpenStudy (callisto):

welcome :)

OpenStudy (callisto):

not really....

OpenStudy (callisto):

Yes :)

OpenStudy (callisto):

It should be :)

OpenStudy (callisto):

welcome :)

OpenStudy (callisto):

cos θ = adj side/ hyp = 8/17 As it is in quad IV, so its value is positive

OpenStudy (callisto):

No problem. Once again, you're not bothering me :)